OpenStudy (anonymous):

Find the value of variables. [-12 -w^2] = [2k -81] [2f 3] [-14 3]

OpenStudy (anonymous):

@iGreen

OpenStudy (gorv):

these are 2 matrics

OpenStudy (gorv):

so when we compare or equate 2 matrices we equate their corresponding element

OpenStudy (gorv):

-12=2k

OpenStudy (gorv):

divide both side by 2 k=-6

OpenStudy (gorv):

so equate similarly equate the element at same places

OpenStudy (anonymous):

So -12=2k k= -6 -w^2=-81(dun know) 2f=-14 f=-7 3=3 =0